OpenStudy (anonymous):

frac{5-x}{x-2}-frac{5x-8}{2-x}

OpenStudy (anonymous):

\[\frac{5-x}{x-2}-\frac{5x-8}{2-x}\]

OpenStudy (anonymous):

Subtract and simplify.

jimthompson5910 (jim_thompson5910):

\[\frac{5-x}{x-2}-\frac{5x-8}{2-x}\] \[\frac{5-x}{x-2}-\frac{5x-8}{-1(x-2)}\] \[\frac{5-x}{x-2}+\frac{5x-8}{x-2}\] \[\frac{5-x+5x-8}{x-2}\] \[\frac{4x-3}{x-2}\]

OpenStudy (anonymous):

Subtract and simplify.

OpenStudy (anonymous):

Thanks again!!

OpenStudy (anonymous):

Subtract and simplify.

OpenStudy (anonymous):

\[f(x)=1/3(x+8)^{2}+1\] Whats the vertex Line of symmetry Maximum/minimum value of f(x)

jimthompson5910 (jim_thompson5910):

Hint: Place in the form \[f(x)=a(x-h)^2+k\] The vertex will be (h,k) and the line of symmetry will be x = h

OpenStudy (anonymous):

would it be (64,1) and -8?

jimthompson5910 (jim_thompson5910):

You have part of it right. You said that the line of symmetry is x = -8. Since h=-8, what does that mean for the vertex?

OpenStudy (anonymous):

-16? or 256?

jimthompson5910 (jim_thompson5910):

Remember, the vertex is (h,k)

OpenStudy (anonymous):

(-8,1)

jimthompson5910 (jim_thompson5910):

Good, you got it. That's the vertex.

OpenStudy (anonymous):

cool, so the line os symmetry would be -8

jimthompson5910 (jim_thompson5910):

x=-8 since this is the equation of the vertical line at -8 on the x-axis

OpenStudy (anonymous):

The line of symmetry cannot be a negative can it?

jimthompson5910 (jim_thompson5910):

it can be any number

OpenStudy (anonymous):

ok would my maximum/minimum be 1 than?

jimthompson5910 (jim_thompson5910):

yes, that is correct. The max/min is 1. The max/min refers to the y coordinate of the vertex.

OpenStudy (anonymous):

Ok thanks again and again, you are a lifesaver!!!